# A unified neural background-error covariance model for midlatitude and tropical atmospheric data assimilation

### v1

@@ 1 Introduction
Recent years have seen a rapid increase in the use of machine learning (ML) to develop novel DA techniques (see Cheng et al. (2023) and Pasmans et al. (2025) for comprehensive reviews). The idea of using ML for unifying the assimilation procedure in the midlatitudes and the tropics was first proposed by Melinc and Zaplotnik (2024, hereafter MZ24), who trained a variational autoencoder (VAE) for reproducing global 850 hPa temperature (T850) fields from ERA5 reanalysis, and performed 3D-Var DA in its reduced-order latent space. Using single observation experiments, they showed that compressing the T850 into a latent space of 100 elements with a VAE produces a background-error covariance matrix $\mathbf{B}_{z}$ that effectively captures local covariances in both midlatitudes and tropics. Notably, although $\mathbf{B}_{z}$ is static and climatological, the background-error covariances in the physical space exhibited state-dependent features. Zheng et al. (2024) then applied the same approach to assimilate sea-surface temperature and demonstrated its feasibility with real-world observations, while Fan et al. (2025) performed 3D-Var and 4D-Var DA in the latent space of a vision transformer. Their analysis increments in the midlatitudes followed the geostrophic balance and were strongly affected by the background flow. Outside of latent-space DA, the physical consistency of analysis increments produced by the ML-based DA techniques received limited attention. Xu et al. (2025) showed that in their system for assimilating satellite observations, an observation with a positive departure of temperature will lead to a physically-consistent decrease in the local relative humidity. Besides, Li et al. (2024) showed physically meaningful analysis increments given the background flow and the time of the observation in their 4D-Var system.
@@ 2 Data and methods
Single observation 3D-Var data assimilation experiments and subsequent forecasts were performed using two neural networks with similar architecture and model variables: (1) a convolutional autoencoder (AE) for performing latent-space 3D-Var (Sec. 2.3), and (2) a convolutional U-Net, based on Perkan et al. (2025), used as a neural-network forecasting model to recursively predict the atmospheric states at 12-hour intervals.
@@ 3.2 Background error-covariance model in the tropics – a response to humidity saturation
The physical balances in the tropics vastly differ from those in the midlatitudes due to a much smaller Coriolis force and tend to be too complex for analytical description. The primary energy source for atmospheric motions in the tropics is the condensation heating in convective cloud systems which drives the large-scale tropical circulation (Holton and Hakim, 2013). The diabatic heating is balanced by the adiabatic cooling of ascending flow (updrafts) inside. This leads to a drop in the surface pressure below the perturbation, establishing a convergent horizontal wind pattern in the lower troposphere that can further fuel the development of a convective system. At the top of the convective system, the statically-stable tropopause acts as a lid for strong vertical motions, leading to the positive pressure perturbation and divergent horizontal outflow. (Gill, 1980; Davey and Gill, 1987).
Although our applied neural networks do not explicitly resolve clouds, condensation, latent heat release, or precipitation, they relate total column water vapour (TCWV) to other atmospheric variables. We tested whether an increase in TCWV in a tropical area with high background TCWV leads to similar effects as described in Davey and Gill (1987) and obtained in simplified numerical studies of tropical response to latent heating (Zaplotnik et al., 2018). Figure 6 shows the analysis increments following an assimilation of TCWV observation in the Central Atlantic Ocean (equator, $33.0\,^{\circ}\mathrm{W}$) with departure $d^{\mathrm{TCWV}}=10$ kg/m2 and error standard deviation $\sigma^{\mathrm{TCWV}}_{o}=3$ kg/m2. The horizontal wind increments at 900 hPa (Fig. 6a) and 200 hPa (Fig. 6b) display convergence in the lower troposphere and divergence in the upper troposphere. At 900 hPa, we also observe a cyclonic vorticity pattern north/south of the perturbation. At 200 hPa, the wind increments are substantially stronger, and the latent-space DA system even generates a Kelvin wave east of the observation. This highlights a remarkable capability of the latent-space background-error covariance model – one not yet demonstrated by any other variational DA system.
@@
Does our DA and the associated latent-space $\mathbf{B}_{z}$-matrix capture the response to latent-heat-induced perturbations in the tropics? To provide further evidence, we analyze the vertical profile of horizontal divergence in the analysis increment. Figure 6e shows that, at the location of the observed TCWV, the convergent wind increment is present in the lower troposphere, peaking around 900 hPa. Near 500 hPa, the flow transitions to divergent, which peaks at 200 hPa, with a magnitude twice as large as the maximum convergence. This structure closely resembles observational findings from Williams and Gray (1973), suggesting that our system realistically represents the dynamical response to tropical latent heating.
@@
Figure 7: Difference between the two forecasts initialized from the analysis and the background, respectively, for a selected ensemble member, based on the experiment in Fig. 6. (a) Difference in the initial condition for TCWV (colors) on top of the 900 hPa background wind (arrows). (b) Difference in the initial condition for the 900 hPa wind (arrows; there is an order of magnitude difference in wind magnitude scaling compared to panel (a)). (c-d) As (a-b), but for the 12-hour forecast lead time. (e-f) As (a-b), but for the 24-hour forecast lead time. (g-h) As (a-b), but for the 48-hour forecast lead time.. The dashed contours in (d,f,h) indicate a negative MSLP difference with 0.1 hPa step, zero contour is omitted.
@@ 4 Estimating latent space background-error covariance model from operational EDA
So far, we have shown that latent-space DA, using a climatological $\mathbf{B}_{z}$-matrix, produces physically reasonable analysis increments both in the tropics and midlatitudes. In this section, we extend our approach by incorporating the ensemble of data assimilation (EDA (Isaksen et al., 2010)), to capture fully flow-dependent background-error variances, commonly referred to as "the errors of the day" (Bonavita et al., 2012).
@@
Climatology-based background-error variances tend to overestimate forecast uncertainty (Bannister, 2008a), whereas ensemble-derived variances often underestimate it. Consequently, ensemble-derived variances are typically smaller than the climatological variability in the AE training set, potentially limiting AE’s ability to reproduce such variances accurately. Additionally, neural networks are known to smooth meteorological fields (Bonavita, 2023), reducing their capacity to reconstruct the smallest-scale features in the ensemble after encoding and decoding. The vast difference in variability is evident in Fig. S6, where rows A and C show that the variances (represented by the standard deviation at each grid point) from the operational ensemble are approximately an order of magnitude smaller than those generated using $\mathbf{B}_{z}^{clim}$ (see Fig. S4, rows C and H). Rows B and D of Fig. S6 display the standard deviation of the 50 ensemble members after processing through the AE. As expected, fine-scale details are lost due to the AE’s limited resolution. Nevertheless, ensemble variability remains reasonably well preserved for the geopotential height fields, which include less variance in the small scales. In contrast, it is poorly reconstructed for other fields – particularly wind components, which have more variance in the small scales.
Using an ensemble of encoded backgrounds and EDA-derived background-error covariance matrix, we repeated a single observation experiment with $d^{\mathrm{Z500}}=30\,\mathrm{m}/\mathrm{s}$ and $\sigma^{\mathrm{Z500}}_{o}=10\,\mathrm{m}/\mathrm{s}$ over Ljubljana. Figure 8a,c-e reveals that the resulting analysis increment preserved the geostrophic and thermal wind balance. To isolate the effect of the $\mathbf{B}_{z}$-matrix, we repeated the experiment once more using the same ensemble of backgrounds, but applied $\mathbf{B}^{clim}_{z}$ for 3D-Var cost function computation (Fig. 8b). In this case, aside from a significantly stronger analysis increment magnitude (approximately five times larger at the observation location) and standard deviation (fourfold increase; not shown), both of which are expected due to an order of magnitude difference in the latent-space variances between the respective $\mathbf{B}_{z}$ matrices, the analysis increments exhibit only minor qualitative differences.
@@
Beyond the differences in the physical space, the analysis increments also differ in the latent space. Figure 9 presents histograms of the relative changes in latent vector elements following 3D-Var cost function minimization across four different experiments. These histograms reveal that, in each experiment, only a small fraction of 12100 elements undergo substantial modification. In the three experiments involving single Z500 observation, the same subset of latent elements is consistently among the most modified, and these changes predominantly affect the region where the analysis increment peaks (Fig. S7). In contrast, these elements are negligibly modified when observing TCWV in the Central Atlantic, where the most adjusted latent element mainly alters the region in the vicinity of that observation (Fig. 9d, Fig. S7, row J).
Figure 9: Histograms of the relative change of respective ensemble members in different experiments. The relative change is calculated as the difference of the latent element’s ensemble mean before and after the assimilation divided by the latent element’s analysis standard deviation. The indices of the most altered element in each of the experiments are highlighted in all four plots. (a) The histogram for the experiment with observing Z500 above Ljubljana using the original setup (Sec. 3.1). (b) The histogram for the experiment with observing Z500 above Ljubljana using the operational ensemble of backgrounds and climatological $\mathbf{B}_{z}$ in 3D-Var cost function minimization. (c) The histogram for the experiment with observing Z500 above Ljubljana using IFS ensemble forecast as the background and its proper $\mathbf{B}_{z}$ in 3D-Var cost function. (d) The histogram for the experiment with observing TCWV above Central Atlantic using the original setup (Sec. 3.2). The bin width is 0.2 on a logarithmic scale.
@@ 5 Discussion, conclusion, and outlook
A key difference between the analysis increments in this study and those presented in MZ24 lies in their horizontal extent – here, the increments are significantly more localized, particularly in the tropics. We found that perturbing individual elements of the latent vector leads to localized changes in the decoded atmospheric fields (Fig. S7). During assimilation, the algorithm primarily adjusted those latent vector elements that describe the features near the observation location (Figs. 9S7). This represents a substantial improvement in the applicability – unlike MZ24, where perturbing individual latent vector elements caused global changes in the decoded fields. This improvement is likely due to the use of a higher-quality autoencoder in the present study, featuring a latent space two orders of magnitude larger than that of MZ24. Incorporating neural networks with even greater representational capacity and larger latent spaces could further enhance localization and provide a more diverse set of learned weather patterns within the latent space.
@@ Acknowledgements
This research was funded by Slovenian Research and Innovation Agency (ARIS) Programme P1-0188 and Grant MR-56969. This research was also supported by the University of Ljubljana Grant SN-ZRD/22-27/0510. Žiga Zaplotnik acknowledges the funding by the European Union under the Destination Earth initiative. The authors are grateful to Nedjeljka Žagar (Universität Hamburg) and Massimo Bonavita (ECMWF) for fruitful discussions on the topic.
@@ A.1 Autoencoder
The structure of the autoencoder (AE), illustrated in Fig. A1, consists of the encoder and the decoder. The encoder $E$ comprises four convolutional blocks, each containing four subblocks with a consistent design. In each block, the input field is first padded using spherical padding, which is periodic in the zonal direction and polar in the meridional direction [Perkan et al., 2025]. The padded field is then passed through a two-dimensional (2D) convolutional layer with a $7\times 7$ kernel and 50 output channels, followed by a leaky rectified linear unit (Leaky ReLU) activation and 2D batch normalization. The stride is set to 1 in all convolutional layers, except in the second subblock of the first block, where it is set to 2. Before entering each subsequent subblock, the output is concatenated with the original input to the first subblock. After each block, the spatial dimensions are reduced using $2\times 2$ max pooling.
@@ Appendix B Appendix B: 3D-Var cost function minimization algorithm
Figure S1: Comparison of the AE outputs with their targets. (A and C) ERA5 reanalysis for the 14th of April, 2020, at 00 UTC. (B and D) Fields from (A and C), after being passed through the AE.
Figure S2: Comparison of the NN forecasts with their targets. (A and C) ERA5 reanalysis for the 15th of April, 2020, at 00 UTC. (B and C) 24-hour forecast, produced by the NN forecasting model, valid at the same time as (A and C).
Figure S3: (a) Comparison of the forecast bias and forecast-error standard deviation for the 24-hour forecast used as the background. The horizontal axis shows the absolute mean error, and the vertical axis shows the standard deviation of the error for each of the 12100 latent vector elements. The statistics were computed over the entire validation set. (b) Distribution of $\mathbf{B}_{z}^{clim}$ elements, shown with a logarithmic bin width of 0.2.
Figure S4: Single observation experiment with observed Z500 above Ljubljana with 30 m departure and 10 m standard deviation. (A and F) Decoded background for all 20 forecasted variables. (B and G) Analysis increments. (C and H) Standard deviation of background. (D and I) Standard deviation of the analysis. (E and J) Ratio between the absolute value of the analysis increment and the analysis standard deviation.
@@
Figure S6: Standard deviation of all variables for the IFS ensemble before (A and C) and after (B and D) propagating it through the autoencoder.
Figure S7: Changes of the decoded fields after modifying a single latent vector element. (A and F) Changes after adding 1 to the latent vector element with index 4786 of the encoded ERA5 reanalysis for the 15th of April, 2020. (B and G) As (A and F), but for the encoded first IFS ensemble member for the 14th of April, 2024. (C and H) As (A and F), but for the latent vector element with index 8878. (D and I) As (A and F), but for the latent vector element with index 9626. (E and J) As (A and F), but for the latent vector element with index 4233. The examined elements are those that are highlighted in Fig. 9.

### v2

@@ 1 Introduction
Recent years have seen a rapid increase in the use of machine learning (ML) to develop novel DA techniques (see Cheng et al. (2023), Bach et al. (2024, Sec. 11.4), and Pasmans et al. (2025) for comprehensive reviews). The idea of using ML for unifying the assimilation procedure in the midlatitudes and the tropics was first proposed by Melinc and Zaplotnik (2024, hereafter MZ24), who trained a variational autoencoder (VAE) for reproducing global 850 hPa temperature (T850) fields from ERA5 reanalysis, and performed 3D-Var DA in its reduced-order latent space. Using single observation experiments, they showed that compressing the T850 into a latent space of 100 elements with a VAE produces a background-error covariance matrix $\mathbf{B}_{z}$ that effectively captures local covariances in both midlatitudes and tropics. Notably, although $\mathbf{B}_{z}$ is static and climatological, the background-error covariances in the physical space exhibited state-dependent features. Zheng et al. (2024) then applied the same approach to assimilate sea-surface temperature and demonstrated its feasibility with real-world observations, while Fan et al. (2025) performed 3D-Var and 4D-Var DA in the latent space of a vision transformer. Their analysis increments in the midlatitudes followed the geostrophic balance and were strongly affected by the background flow. Outside of latent-space DA, the physical consistency of analysis increments produced by the ML-based DA techniques received limited attention. Xu et al. (2025) showed that in their system for assimilating satellite observations, an observation with a positive departure of temperature will lead to a physically-consistent decrease in the local relative humidity. Besides, Li et al. (2024) showed physically meaningful analysis increments given the background flow and the time of the observation in their 4D-Var system.
@@ 2 Data and methods
Single observation 3D-Var data assimilation experiments and subsequent forecasts were performed using two neural networks with similar architecture and model variables: (1) a convolutional autoencoder (AE) for performing latent-space 3D-Var (Sec. 2.3), and (2) a convolutional U-Net, based on Perkan and Zaplotnik (2025), used as a neural-network forecasting model to recursively predict the atmospheric states at 12-hour intervals.
@@ 3.2 Background error-covariance model in the tropics – a response to humidity saturation
The physical balances in the tropics vastly differ from those in the midlatitudes due to a much smaller Coriolis force and tend to be too complex for analytical description. The primary energy source for atmospheric motions in the tropics is the condensation heating in convective cloud systems which drives the large-scale tropical circulation (Holton and Hakim, 2013). The diabatic heating is balanced by the adiabatic cooling of ascending flow (updrafts) inside. This leads to a drop in the surface pressure below the perturbation, establishing a convergent horizontal wind pattern in the lower troposphere that can further fuel the development of a convective system. At the top of the convective system, the statically-stable tropopause acts as a lid for strong vertical motions, leading to the positive pressure perturbation and divergent horizontal outflow (Gill, 1980; Davey and Gill, 1987).
Although our applied neural networks do not explicitly resolve clouds, condensation, latent heat release, or precipitation, they relate total column water vapour (TCWV) to other atmospheric variables. We tested whether an increase in TCWV in a tropical area with high background TCWV leads to similar effects as described in Davey and Gill (1987) and obtained in simplified numerical studies of tropical response to latent heating (Zaplotnik et al., 2018). Figure 6 shows the analysis increments following an assimilation of TCWV observation in the Central Atlantic Ocean (equator, $33.0\,^{\circ}\mathrm{W}$) with departure $d^{\mathrm{TCWV}}=10$ kg/m2 and error standard deviation $\sigma^{\mathrm{TCWV}}_{o}=3$ kg/m2. The horizontal wind increments at 900 hPa (Fig. 6a) and 200 hPa (Fig. 6b) display convergence in the lower troposphere and divergence in the upper troposphere. At 900 hPa, we also observe a cyclonic vorticity pattern north/south of the perturbation. At 200 hPa, the wind increments are substantially stronger, and the latent-space DA system even generates a Kelvin wave east of the observation. The analysis increments that span the whole tropospheric column are a consequence of tropospheric-wide background-error covariances that are typical of precipitating regions where condensational latent heating occurs Li et al. (2023).
@@
Does our DA and the associated latent-space $\mathbf{B}_{z}$-matrix capture the response to latent-heat-induced perturbations in the tropics? To provide further evidence, we analyze the vertical profile of horizontal divergence in the analysis increment. Figure 6e shows that, at the location of the observed TCWV, the convergent wind increment is present in the lower troposphere, peaking around 900 hPa. Near 500 hPa, the flow transitions to divergent, which peaks at 200 hPa, with a magnitude twice as large as the maximum convergence. This vertical structure closely resembles observational findings from Williams and Gray (1973), suggesting that our system realistically represents the dynamical response to tropical latent heating. This highlights a remarkable capability of the latent-space background-error covariance model – one not yet demonstrated by any other variational DA system.
@@
Figure 7: Difference between the two forecasts initialized from the analysis and the background from the single-observation experiment in Fig. 6. (a) Difference in the initial condition for TCWV (colors) on top of the 900 hPa background wind (arrows). (b) Difference in the initial condition for the 900 hPa wind (arrows; there is an order of magnitude difference in wind magnitude scaling compared to panel (a)). (c-d) As (a-b), but for the 12-hour forecast lead time. (e-f) As (a-b), but for the 24-hour forecast lead time. (g-h) As (a-b), but for the 48-hour forecast lead time.. The dashed contours in (d,f,h) indicate a negative MSLP difference with 0.1 hPa step, zero contour is omitted.
@@ 4 Estimating latent space background-error covariance model from operational EDA
So far, we have shown that latent-space DA, using a climatological $\mathbf{B}_{z}$-matrix, produces physically reasonable analysis increments both in the tropics and midlatitudes. In this section, we extend our approach by incorporating the ensemble of data assimilation (EDA, (Isaksen et al., 2010)) to capture fully flow-dependent background-error variances, commonly referred to as "the errors of the day" (Bonavita et al., 2012).
@@
Climatology-based background-error variances tend to overestimate forecast uncertainty (Bannister, 2008a), whereas ensemble-derived variances often underestimate it. The vast difference in variances is evident in Fig. S6, which shows that the variances from the operational ensemble are approximately an order of magnitude smaller than those in $\mathbf{B}_{z}^{clim}$ (compare to Fig. S4, rows C and H). Moreover, neural networks tend to smooth meteorological fields (Bonavita, 2023), reducing their ability to reconstruct the fine-scale features after encoding and decoding. For example, Fig. S6 displays how the standard deviation of the 50 ensemble members changes after autoencoder processing, with noticeable loss of fine-scale details due to the AE’s limited resolution. While ensemble variance is reasonably preserved for the geopotential height, it is poorly reconstructed for fields like winds and specific humidity, which contain greater small-scale variance.
Using an ensemble of encoded backgrounds from EDA and EDA-derived background-error covariance matrix (Eq. 7), we repeated a single-observation experiment with $d^{\mathrm{Z500}}=30\,\mathrm{m}/\mathrm{s}$ and $\sigma^{\mathrm{Z500}}_{o}=10\,\mathrm{m}/\mathrm{s}$ over Ljubljana. Figure 8a,c-e reveals that the resulting analysis increment preserved the geostrophic and thermal wind balance. To isolate the effect of the $\mathbf{B}_{z}$-matrix from the impact of using operational ensemble of backgrounds, we repeated the experiment once more using the same ensemble of backgrounds, but applied $\mathbf{B}^{clim}_{z}$ for 3D-Var cost function computation (Fig. 8b). Qualitatively, the analysis increments exhibited only minor differences. However, the magnitude of the analysis increment was significantly larger using $\mathbf{B}^{clim}_{z}$ (approximately five times larger at the observation location), while the background-error standard deviation was larger too (fourfold increase; not shown), both of which are expected due to an order of magnitude difference in the latent-space variances between the respective $\mathbf{B}_{z}$ matrices.
@@
Beyond the differences in the physical space, the analysis increments also differ in the latent space. Figure 9 presents histograms of the relative changes in latent vector elements (i.e., ratio of the analysis increment and standard deviation) following 3D-Var cost function minimization across four different experiments. These histograms reveal that, in each experiment, only a small fraction of 12100 elements undergo substantial modification. In the three experiments involving single Z500 observation, the same subset of latent elements is consistently among the most modified, and these changes predominantly affect the region where the analysis increment peaks (Fig. S7). In contrast, these elements are negligibly modified when observing TCWV in the Central Atlantic, where the most adjusted latent element mainly alters the region in the vicinity of that observation (Fig. 9d, Fig. S7, row J).
Figure 9: Histograms of the relative changes of latent vector elements in different single-observation experiments. The indices of the most altered element in each of the experiments are highlighted in all four plots. (a) The histogram for the experiment with observing Z500 above Ljubljana using the original setup (Sec. 3.1). (b) The histogram for the experiment with observing Z500 above Ljubljana using the operational ensemble of backgrounds and climatological $\mathbf{B}_{z}$ in 3D-Var cost function minimization. (c) The histogram for the experiment with observing Z500 above Ljubljana using IFS ensemble forecast as the background and its proper $\mathbf{B}_{z}$ in 3D-Var cost function. (d) The histogram for the experiment with observing TCWV above Central Atlantic using the original setup (Sec. 3.2). The bin width is 0.2 on a logarithmic scale.
@@ 5 Discussion, conclusion, and outlook
A key difference between the analysis increments in this study and those presented in MZ24 lies in their horizontal extent – here, the increments are significantly more localized, particularly in the tropics. We found that perturbing individual elements of the latent vector leads to localized changes in the decoded atmospheric fields (Fig. S7). During assimilation, the algorithm primarily adjusted those latent vector elements that describe the features near the observation location (Figs. 9, S7). This represents a substantial improvement in the applicability – unlike MZ24, where perturbing individual latent vector elements caused global changes in the decoded fields. This improvement is likely due to the use of a higher-quality autoencoder in the present study, featuring a latent space two orders of magnitude larger than that of MZ24. Incorporating neural networks with even greater representational capacity and larger latent spaces could further enhance localization and provide a more diverse set of learned weather patterns within the latent space.
@@ Acknowledgements
This research was funded by the Slovenian Research and Innovation Agency (ARIS) Programme P1-0188, Grant MR 56969, and Grant 0510–1554-58115. This research was also supported by the University of Ljubljana Grant SN-ZRD/22-27/0510. Žiga Zaplotnik acknowledges the funding by the European Union under the Destination Earth initiative. The authors are grateful to Nedjeljka Žagar (Universität Hamburg) and Massimo Bonavita (ECMWF) for fruitful discussions on the topic.
@@ A.1 Autoencoder
The structure of the autoencoder (AE), illustrated in Fig. A1, consists of the encoder and the decoder. The encoder $E$ comprises four convolutional blocks, each containing four subblocks with a consistent design. In each block, the input field is first padded using spherical padding, which is periodic in the zonal direction and polar in the meridional direction [Perkan and Zaplotnik, 2025]. The padded field is then passed through a two-dimensional (2D) convolutional layer with a $7\times 7$ kernel and 50 output channels, followed by a leaky rectified linear unit (Leaky ReLU) activation and 2D batch normalization. The stride is set to 1 in all convolutional layers, except in the second subblock of the first block, where it is set to 2. Before entering each subsequent subblock, the output is concatenated with the original input to the first subblock. After each block, the spatial dimensions are reduced using $2\times 2$ max pooling.
@@ Appendix B Appendix B: 3D-Var cost function minimization algorithm
Figure S1: Comparison of the autoencoder (AE) outputs with their ERA5 targets. (Rows A, C) ERA5 reanalysis on April 14, 2020, at 00 UTC. (B, D) Corresponding AE reconstructions of the fields shown in (A, C).
Figure S2: Comparison of the neural network (NN) forecasts with ERA5 reanalysis. (Rows A, C) ERA5 fields on April 15, 2020, at 00 UTC. (B, D) 24-hour forecasts from the NN model, valid at the same time.
Figure S3: (a) Comparison of the forecast bias and forecast-error standard deviation for the 24-hour forecast used as background. Each point represents one of the 12100 latent vector elements, with absolute mean error on the x-axis, and error standard deviation on the y-axis, computed over the entire validation set. (b) Distribution of $\mathbf{B}_{z}^{clim}$ elements, shown using logarithmic bin width of 0.2.
Figure S4: Single-observation DA experiment using a Z500 observation over Ljubljana, Slovenia, with a 30 m departure and a 10 m observation-error standard deviation. (Rows A and F) Decoded background fields for all 20 forecasted variables. (B and G) Analysis increments. (C and H) Background-error standard deviation. (D and I) Analysis-error standard deviation. (E and J) Ratio of the absolute value of the analysis increment and the analysis-error standard deviation.
@@
Figure S6: Standard deviation of all variables for the IFS ensemble before (rows A and C) and after (B and D) propagating it through the autoencoder.
Figure S7: Changes of the decoded fields after modifying a single latent vector element. (Rows A and F) Changes after adding 1 to the latent vector element with index 4786 of the encoded ERA5 reanalysis for the 15th of April, 2020. (B and G) As (A and F), but for the encoded first IFS ensemble member for the 14th of April, 2024. (C and H) As (A and F), but for the latent vector element with index 8878. (D and I) As (A and F), but for the latent vector element with index 9626. (E and J) As (A and F), but for the latent vector element with index 4233. The examined elements are those that are highlighted in Fig. 9.